OpenStudy (anonymous):

Please help Solve x2 − 12x + 5 = 0 using the completing-the-square method. x = six plus or minus the square root of five x = negative six plus or minus the square root of five x = six plus or minus the square root of thirty one x = negative six plus or minus the square root of thirty one

OpenStudy (anonymous):

The answers should look like https://gyazo.com/872a0909fd0d54eba579498515313f6f the one with the answer is my best guess

OpenStudy (jadedry):

\[x^2 - 12x + (12/2)^2 = -5 + (12/2)^2\] = \[x^2 -12x + 36 = 31\] = \[(x-6)^2 = 31\] =\[x-6 = \pm \sqrt31 \] therefore \[x = 6 \pm \sqrt 31\]

OpenStudy (anonymous):

Oh thank you very much!

OpenStudy (jadedry):

@AntiFox No problem! c: